OpenStudy (anonymous):

Find the polar coordinates (-4√(3),-4) for r > 0

OpenStudy (anonymous):

\[x = rcos \theta, y = rsin \theta\] \[4\sqrt{3} = rcos \theta\] \[-4=rsin \theta\] From that point you can get theta. Then... think a bit more to get r.

OpenStudy (anonymous):

So rsin theta equals -4?

OpenStudy (anonymous):

Yes, the you need to do: -rsin theta = 4 = (rcos theta)/sqrt(3)

OpenStudy (anonymous):

tan theta equals....